OpenStudy (anonymous):

At What points on the graph of f(x)=18x-2x^3 is the slope equal to -6

OpenStudy (anonymous):

slope is given by df(x)/dx = 18 - 6x^2 in this case. slope = -6 implies, 18 - 6x^2 = -6 Therefore, 6x^2 = 24 Therefore, x^2 = 4 Therefore, x = +sqrt(4) or - sqrt(4) = +2 or -2 Therefore, at x = 2 and x = -2, the slope = -6